A car is moving on a horizontal road with a speed of 20 m/s. Rain is falling vertically downwards with a speed of 10m/s. If the raindrops strike the windscreen of the car normally, then the angle made by the windscreen with the vertical?


Ans tan-1(.5)

Grade:

1 Answers

Vanya Saxena
14 years ago

Let the velocity of the car be Vc and that of the rain be Vr Now on plotting the figure we get a right angled triangle with its perpendicular as Vr=10m/s and the horizontal component i.e. the base as Vc=20 m/s Now since the angle required is the vertical angle off the triangle

therefore

                                tanΦ=Vc/Vr

                                tan Φ=1/2

                                   Φ=tan-1(.5)